# Cuspidal edges on focal surfaces of regular surfaces with parabolic points

### v1

@@ Abstract.
parallel surface, focal surface, singularity, cuspidal lips, cuspidal beaks, Gaussian curvature
@@ 1. Introduction
For a regular surface in Euclidean $3$-space $\mathbb{R}^{3}$, its parallels and focal surfaces can be examined. While the initial surface is non-singular, its parallels and focal surfaces generally have singularities ([10]). From the viewpoint of singularity theory of smooth mappings, singularities that appear on parallel surfaces are Legendrian singularities, while those on focal surfaces are Lagrangian singularities ([1, 2, 10]). In both cases, a cuspidal edge is known to be a generic singularity ([1]). It has been established that there is a relationship between singularities of parallels and focal surfaces of a regular surface ([8, 26]). See Table 1 for example. We remark that parallel surfaces are fronts ([1, 2, 10, 28]).
@@
Table 1. Relation between singularities (see Proposition 4.1).
Conversely, there are numerous studies on surface with cuspidal edge singularities from the perspective of differential geometry (cf. [7, 16, 14, 19, 22, 24, 25]). At a cuspidal edge, several geometric invariants are introduced ([16, 14, 22]). In particular, geometric invariants called the limiting normal curvature and the singular curvature play a crucial role. It has been established that the limiting normal curvature is associated with the boundedness of the Gaussian curvature at a cuspidal edge ([22, 16]). While the singular curvature relates to convexity or concavity of cuspidal edge. More precisely, the following fact is known ([22, Theorem 3.1]): The Gaussian curvature is bounded near a cuspidal edge if and only if the limiting normal curvature vanishes along the edge. Moreover, if the Gaussian curvature takes non-negative values near a cuspidal edge, then the singular curvature is non-positive at the cuspidal edge. For a pseudo-spherical front, that is, a surface with certain singularities whose Gaussian curvature is negative constant on the set of regular points, the Gaussian curvature of its focal surface is bounded (cf. (A.7)). In addition, we see the following assertions ([27]): If the initial pseudo-spherical front has a cuspidal lips (resp. cuspidal beaks), then the singular curvature of the focal surface is positive (resp. negative) at the corresponding cuspidal edge. Thus it is natural to ask that how singularities of fronts (parallel surfaces) relate to geometric properties of cuspidal edges on focal surfaces.
In this paper, we investigate the limiting normal curvature and the singular curvature at cuspidal edges of focal surfaces of regular surfaces. In particular, we deal with a surface whose Gauss map has singularity at some point. In that case, the Gaussian curvature of parallel surfaces satisfies certain bounded condition (see Propositions 3.3 and 3.4). Moreover, the limiting normal curvature defined at a corank one singular point on the parallel surface vanishes in the case (Proposition 3.5). Comparing conditions for singularities of parallel surfaces and focal surfaces, it holds that the focal surface is a cuspidal edge if the parallel surface is a swallowtail, a cuspidal edge or a cuspidal beaks (Proposition 4.1). For a curvature line parametrized surfaces, we give explicit representations for the limiting normal curvature and the singular curvature at a cuspidal edge of the focal surface (Propositions 4.3 and 4.5). In particular for the singular curvature, we shall show that relation between the sign of the singular curvature at a cuspidal edge of the focal surface and singularities of parallel surface (Theorem 4.6).
@@ Acknowledgments
The author thanks Kentaro Saji and Runa Shimada for fruitful discussions and valuable comments.
@@ 2.1. Surfaces in Euclidean $3$-space
We here review some fundamental concepts from differential geometry of immersed surfaces in Euclidean $3$-space $\mathbb{R}^{3}$. For details, see [6, 9, 21].
2.1.1. Surfaces parametrized by curvature line coordinates
Let $U\subset\mathbb{R}^{2}$ be a domain. Let $f\colon U\to\mathbb{R}^{3}$ be an immersion and $\nu$ its unit normal vector. Then we call the image of $f$ or $f$ itself a surface. We assume that $f$ is a surface without umbilic point in the following. Then we can take a curvature line coordinate system $(u,v)$ on $U$. In this case, it holds that
$\left\langle{f_{u}},{f_{v}}\right\rangle=\left\langle{\nu_{u}},{f_{v}}\right\rangle=\left\langle{\nu_{v}},{f_{u}}\right\rangle=0,$
where $\left\langle{\cdot},{\cdot}\right\rangle$ is the canonical inner product on $\mathbb{R}^{3}$, $(\cdot)_{u}=\partial(\cdot)/\partial u$ and $(\cdot)_{v}=\partial(\cdot)/\partial v$. Moreover, the first and second fundamental forms of $f$ can be written as
$\operatorname{I}=Edu^{2}+Gdv^{2},\quad\operatorname{II}=Ldu^{2}+Ndv^{2},$
where
$E=\left\langle{f_{u}},{f_{u}}\right\rangle,\quad G=\left\langle{f_{v}},{f_{v}}\right\rangle,\quad L=-\left\langle{f_{u}},{\nu_{u}}\right\rangle,\quad N=-\left\langle{f_{v}},{\nu_{v}}\right\rangle.$
Then the Gauss and Weingarten formulas are as follows:
@@ 2.1.1. Surfaces parametrized by curvature line coordinates
where $\kappa_{1}$ and $\kappa_{2}$ are principal curvatures of $f$ given by
$\kappa_{1}=\frac{L}{E},\quad\kappa_{2}=\frac{N}{G}.$
Moreover, Codazzi equations can be written as
@@
Definition 2.2.
For a surface $f\colon U\to\mathbb{R}^{3}$ parametrized by curvature line coordinate system $(u,v)$,
  • a point $p\in U$ is said to be a $\partial_{u}$-ridge point (resp. $\partial_{v}$-ridge point) if $(\kappa_{1})_{u}(p)=0$ (resp. $(\kappa_{2})_{v}(p)=0$);
  • a point $p\in U$ is said to be a $\partial_{u}$-sub-parabolic point (resp. $\partial_{v}$-sub-parabolic point) if $(\kappa_{2})_{u}(p)=0$ (resp. $(\kappa_{1})_{v}(p)=0$).
Ridge points on a surface are originally introduced by Porteous to investigate singularities of distance squared functions on the surface ([20]). For details of ridge and sub-parabolic points on a surface, see [4, 3, 20, 21, 18].
2.1.2. Curvature line frame
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface without umbilic point. Let $\nu$ be a unit normal vector to $f$. We then take a curvature line coordinate system $(u,v)$ on $U$. We define maps $\bm{e}_{i}\colon U\to\mathbb{R}^{3}$ ($i=1,2$) by
(2.3) $\bm{e}_{1}=\frac{f_{u}}{|f_{u}|},\quad\bm{e}_{2}=\frac{f_{v}}{|f_{v}|}.$
Then it holds that
$\left\langle{\bm{e}_{i}},{\bm{e}_{j}}\right\rangle=\delta_{ij},\quad\nu=\bm{e}_{1}\times\bm{e}_{2},$
where $\delta_{ij}$ is the Kronecker delta. Thus the triplet $\{\bm{e}_{1},\bm{e}_{2},\nu\}$ forms an orthonormal frame along $f$. We call this frame the curvature line frame along $f$.
Lemma 2.3 (cf. [5]).
With the above condition, we have
$\begin{pmatrix}\bm{e}_{1}\\ \bm{e}_{2}\\ \nu\end{pmatrix}_{u}=\begin{pmatrix}0&x_{1}&x_{2}\\ -x_{1}&0&0\\ -x_{2}&0&0\end{pmatrix}\begin{pmatrix}\bm{e}_{1}\\ \bm{e}_{2}\\ \nu\end{pmatrix},\quad\begin{pmatrix}\bm{e}_{1}\\ \bm{e}_{2}\\ \nu\end{pmatrix}_{v}=\begin{pmatrix}0&y_{1}&0\\ -y_{1}&0&y_{3}\\ 0&-y_{3}&0\end{pmatrix}\begin{pmatrix}\bm{e}_{1}\\ \bm{e}_{2}\\ \nu\end{pmatrix}$,
where
$x_{1}=\frac{(\kappa_{1})_{v}|f_{u}|}{(\kappa_{1}-\kappa_{2})|f_{v}|},\quad x_{2}=\kappa_{1}|f_{u}|,\quad y_{1}=\frac{(\kappa_{2})_{u}|f_{v}|}{(\kappa_{1}-\kappa_{2})|f_{u}|},\quad y_{3}=\kappa_{2}|f_{v}|.$
Proof.
For $\nu$, it holds that
$\nu_{u}=-\kappa_{1}f_{u}=-\kappa_{1}|f_{u}|\bm{e}_{1},\quad\nu_{v}=-\kappa_{2}f_{v}=-\kappa_{2}|f_{v}|\bm{e}_{2}$
by Weingarten formulas (2.1). By (2.3) and Lemma 2.1, $(\bm{e}_{1})_{u}$ and $(\bm{e}_{2})_{u}$ can be calculated as
$(\bm{e}_{1})_{u}$ $=\frac{f_{uu}}{|f_{u}|}-\frac{f_{u}(|f_{u}|)_{u}}{|f_{u}|^{2}}=\frac{1}{|f_{u}|}\left(\Gamma^{1}_{11}f_{u}+\Gamma^{2}_{11}f_{v}+L\nu\right)-\Gamma^{1}_{11}\bm{e}_{1}$
$=-\frac{E_{v}}{2G|f_{u}|}f_{v}+\frac{L}{|f_{u}|}\nu=\frac{(\kappa_{1})_{v}|f_{u}|}{(\kappa_{1}-\kappa_{2})|f_{v}|}\bm{e}_{2}+\kappa_{1}|f_{u}|\nu,$
$(\bm{e}_{2})_{u}$ $=\frac{f_{uv}}{|f_{v}|}-\frac{f_{v}(|f_{v}|)_{u}}{|f_{v}|^{2}}=\frac{1}{|f_{v}|}\left(\Gamma^{1}_{12}f_{u}+\Gamma^{2}_{12}f_{v}\right)-\Gamma^{2}_{12}\bm{e}_{2}$
$=\frac{E_{v}}{2E|f_{v}|}f_{u}=-\frac{(\kappa_{1})_{v}}{(\kappa_{1}-\kappa_{2})|f_{v}|}f_{u}=-\frac{(\kappa_{1})_{v}|f_{u}|}{(\kappa_{1}-\kappa_{2})|f_{v}|}\bm{e}_{1}.$
By similar calculations, we get
$(\bm{e}_{1})_{v}$ $=\frac{f_{uv}}{|f_{u}|}-\frac{\left\langle{f_{u}},{f_{uv}}\right\rangle}{|f_{u}|^{2}}=\frac{\Gamma^{2}_{12}}{|f_{u}|}f_{v}=\frac{G_{u}}{2G|f_{u}|}f_{v}=\frac{(\kappa_{2})_{u}|f_{v}|}{(\kappa_{1}-\kappa_{2})|f_{u}|}\bm{e}_{2},$
$(\bm{e}_{2})_{v}$ $=\frac{f_{vv}}{|f_{v}|}-\Gamma^{2}_{22}\bm{e}_{2}=\frac{1}{|f_{v}|}\left(\Gamma^{1}_{22}f_{u}+N\nu\right)=-\frac{G_{u}}{2E|f_{v}|}f_{u}+\frac{N}{|f_{v}|}\nu$
$=-\frac{(\kappa_{2})_{u}G}{(\kappa_{1}-\kappa_{2})E|f_{v}|}f_{u}+\kappa_{2}|f_{v}|\nu=-\frac{(\kappa_{2})_{u}|f_{v}|}{(\kappa_{1}-\kappa_{2})|f_{u}|}\bm{e}_{1}+\kappa_{2}|f_{v}|\nu.$
Thus we have the assertions. ∎
By this lemma, it holds that
$(\bm{e}_{1})_{u}\times(\bm{e}_{1})_{v}=-\frac{\kappa_{1}(\kappa_{2})_{u}|f_{v}|}{\kappa_{1}-\kappa_{2}}\bm{e}_{1},\quad(\bm{e}_{2})_{u}\times(\bm{e}_{2})_{v}=-\frac{\kappa_{2}(\kappa_{1})_{v}|f_{u}|}{\kappa_{1}-\kappa_{2}}\bm{e}_{2}.$
Therefore sets of singularities of $\bm{e}_{1}$ and $\bm{e}_{2}$ are
(2.4) $S(\bm{e}_{1})$ $=\{q\in U\mid\kappa_{1}(q)=0\}\cup\{q\in U\mid(\kappa_{2})_{u}(q)=0\},$
$S(\bm{e}_{2})$ $=\{q\in U\mid\kappa_{2}(q)=0\}\cup\{q\in U\mid(\kappa_{1})_{v}(q)=0\},$
respectively. Since $S(\nu)=\kappa_{1}^{-1}(0)\cup\kappa_{2}^{-1}(0)$, if $p\in S(\nu)$, then $p\in S(\bm{e}_{1})$ or $p\in S(\bm{e}_{2})$. See [5] for singularities of frame mappings on regular surfaces.
@@ 2.2. Fronts
To investigate parallel surfaces below, we here recall the class of surface with certain singularities called fronts ([1, 28]). Let $f\colon U\to\mathbb{R}^{3}$ be a $C^{\infty}$ map, where $U\subset\mathbb{R}^{2}$ is an open set. Then $f$ is said to be a front if there exists a $C^{\infty}$ map $\nu\colon U\to\mathbb{S}^{2}$ such that
@@
Let $f\colon U\to\mathbb{R}^{3}$ be a front with a unit normal vector field $\nu$. Then a point $p\in U$ is said to be a singular point of $f$ if $f$ is not an immersion at $p$, that is, $\operatorname{rank}df_{p}<2$. We denote by $S(f)$ the set of singular points of $f$. We define a function $\lambda\colon U\to\mathbb{R}$ by
@@
We then call $\lambda$ the signed area density function of $f$. It is clear that $S(f)=\lambda^{-1}(0)$ holds. Moreover, a function $\tilde{\lambda}$ which is non-zero proportional to $\lambda$ is called an identifier of singularities of $f$. A singular point $p\in S(f)$ of $f$ is said to be of non-degenerate if $(\tilde{\lambda}_{u},\tilde{\lambda}_{v})(p)\neq(0,0)$. Otherwise, it is said to be of degenerate. For a non-degenerate singular point $p\in S(f)$, there exist a neighborhood $V(\subset U)$ of $p$ and a regular curve $\gamma\colon(-\varepsilon,\varepsilon)\to V$ with $\gamma(0)=p$ such that $\tilde{\lambda}(\gamma(t))=0$ for $t\in(-\varepsilon,\varepsilon)$. This curve $\gamma$ is called a singular curve for $f$ passing through $p=\gamma(0)$. On the other hand, for a singular point $p\in S(f)$ satisfying $\operatorname{rank}df_{p}=1$, there exists a non-zero vector field $\eta$ on $U$ such that $\eta_{p}$ generates the kernel of $df_{p}$. This vector field $\eta$ is called a null vector field for $f$.
@@
Let $f_{i}\colon(U_{i},p_{i})\to\mathbb{R}^{3}$ ($i=1,2$) be $C^{\infty}$ map germs. Then $f_{1}$ and $f_{2}$ are $\cal{A}$-equivalent if there exist diffeomorphism germs $\varphi\colon(U_{1},p_{1})\to(U_{2},p_{2})$ and $\Phi\colon(\mathbb{R}^{3},f_{1}(p_{1}))\to(\mathbb{R}^{3},f_{2}(p_{2}))$ such that $\Phi\circ f_{1}=f_{2}\circ\varphi$.
@@
    $f$ is a cuspidal edge at $p(\in U)$ if $f$ is $\cal{A}$-equivalent to the germ $(u,v)\mapsto(u,v^{2},v^{3})$ at the origin;
@@
    $f$ is a swallowtail at $p$ if $f$ is $\cal{A}$-equivalent to the germ $(u,v)\mapsto(u,4v^{3}+2uv,3v^{4}+uv^{2})$ at the origin;
@@
    $f$ is a cuspidal butterfly at $p$ if $f$ is $\cal{A}$-equivalent to the germ $(u,v)\mapsto(u,5v^{4}+2uv,4v^{5}+uv^{2})$ at the origin;
@@
    $f$ is a cuspidal lips at $p$ if $f$ is $\cal{A}$-equivalent to the germ $(u,v)\mapsto(u,2v^{3}+u^{2}v,3v^{4}+u^{2}v^{2})$ at the origin;
@@
    $f$ is a cuspidal beaks at $p$ if $f$ is $\cal{A}$-equivalent to the germ $(u,v)\mapsto(u,2v^{3}-u^{2}v,3v^{4}-u^{2}v^{2})$ at the origin.

### v2

@@ Abstract.
parallel surface, focal surface, cuspidal edge
@@ 1. Introduction
Let $f\colon U\to\mathbb{R}^{3}$ be a $C^{\infty}$ regular surface, where $U$ is an open set in $\mathbb{R}^{2}$. Then for a fixed $t\in\mathbb{R}$, a parallel surface $f^{t}\colon U\to\mathbb{R}^{3}$ of $f$ is defined by
$f^{t}(u,v)=f(u,v)+t\nu(u,v)\quad((u,v)\in U),$
where $\nu$ is the unit normal vector field along $f$ given by $\nu=(f_{u}\times f_{v})/|f_{u}\times f_{v}|$. A parallel surface is a (wave) front in $\mathbb{R}^{3}$ ([1, 2, 25]). Let $\kappa_{i}$ ($i=1,2$) be principal curvatures of $f$. Then $f^{t}$ is singular at $p\in U$ when $t=1/\kappa_{i}(p)$ ($i=1$ or $2$). Singularities of parallel surfaces are Legendrian singularities. If $p$ is not an umbilic point of $f$, that is, $\kappa_{1}(p)\neq\kappa_{2}(p)$, then $f^{t}$ may be a cuspidal edge, a swallowtail, a cuspidal lips, a cuspidal beaks and a cuspidal butterfly at $p$ (see Definition 2.3; cf. [1, Chapters 21 and 22]). Characterizations for these singularities are given by Fukui and Hasegawa [5, Theorems 3.4 and 3.5] in terms of geometric properties of the initial surface. On the other hand, one can define focal surfaces $C_{i}\colon U\setminus\kappa_{i}^{-1}(0)\to\mathbb{R}^{3}$ ($i=1,2$) associated with $\kappa_{i}$ by
$C_{i}(u,v)=f(u,v)+\frac{1}{\kappa_{i}(u,v)}\nu(u,v).$
The focal surface $C_{i}$ corresponds to the locus of centers of principal curvature spheres with respect to $\kappa_{i}$. The relationship between the Gaussian curvature of the focal surface and geometric properties of the initial surface is well understood (see [4, page 183]). Moreover, it is known that focal surfaces have Lagrangian singularities ([1, Chapter 21]). Furthermore, it can be seen that the singularities of parallel surfaces are located on a focal surface. The relationships between several singularities of parallel surfaces and of focal surfaces are shown in Table 1 (see Proposition 4.2).
Table 1. Relation between singularities.
@@
On the other hand, there are various studies on surface with cuspidal edge singularities from differential geometric viewpoint (cf. [6, 12, 13, 14, 16, 20, 21, 22, 24]). At a cuspidal edge, several geometric invariants are introduced ([12, 13, 20]). The limiting normal curvature $\kappa_{\nu}$ and singular curvature $\kappa_{s}$ are known as principal geometric invariants at a cuspidal edge. The limiting normal curvature is related to the behavior of the Gaussian curvature near a cuspidal edge ([12, Theorem A]; see Figure 1). (We remark that the limiting normal curvature can be defied at a rank one singular point of a front ([12, pages 253–254]).)
(a) $\kappa_{\nu}=0$
(b) $\kappa_{\nu}\neq 0$
Figure 1. Cuspidal edges with vanishing $\kappa_{\nu}$ (left) and non-vanishing $\kappa_{\nu}$ (right). Arrows indicate unit normal vectors.
Conversely, the sign of the singular curvature reflects the concavity or convexity of the cuspidal edge (see Figure 2; [20, Corollary 1.18]).
(a) $\kappa_{s}>0$
(b) $\kappa_{s}<0$
Therefore examining the singular curvature at a cuspidal edge on focal surfaces determines its concavity or convexity. Consequently, by investigating the properties of the initial surface or parallel surface that determines the sign of the singular curvature at cuspidal edges appearing on the focal surface, it is expected that the geometric properties of the focal surface will become clearer.
In this paper, we study geometric properties of parallel surfaces and focal surfaces at rank one singular points. In particular, we focus on relationships between geometric perspective and behavior of the limiting normal and singular curvatures. In Section 3, we investigate parallel surfaces. When the initial surface does not have umbilic points, its parallel surfaces may have rank one singularities. Thus one can consider the limiting normal curvature at that point on a parallel surface. In particular, the following results hold.
Proposition 1.1.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface without umbilic point and $\nu$ its unit normal vector. Suppose that the principal curvature $\kappa_{1}$ (resp. $\kappa_{2}$) does not vanish on $U$. If the parallel surface $f^{t}$ with $t=1/\kappa_{1}(p)$ (resp. $t=1/\kappa_{2}(p)$) satisfies $\operatorname{rank}df^{t}_{p}=1$, then the limiting normal curvature $\kappa_{\nu}^{t}$ of $f^{t}$ satisfies
$\kappa_{\nu}^{t}=\frac{\kappa_{1}\kappa_{2}}{\kappa_{1}-\kappa_{2}}\quad\left(\textit{resp.}~\kappa_{\nu}^{t}=\frac{\kappa_{1}\kappa_{2}}{\kappa_{2}-\kappa_{1}}\right)$
at $p$.
In Section 4, we deal with focal surfaces. Especially, we study the limiting normal curvature and the singular curvature at a cuspidal edge on the focal surface. The curvature line coordinate system is employed to provide explicit representations of these curvatures. As a consequence, we establish a geometric criterion that determines the sign of the singular curvature of the focal surface in terms of the type of singularity of the parallel surface. More precisely, we shall show the following.
Theorem 1.2.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface whose parallel surface $f^{t}$ is a cuspidal lips (resp. a cuspidal beaks) at $p\in U$, where $t=1/\kappa_{i}(p)$ ($i=1$ or $2$). Suppose that the limiting normal curvature $\kappa_{\nu}^{t}$ of $f^{t}$ vanishes at $p$. Then the singular curvature $\kappa_{s}^{C_{i}}$ at a cuspidal edge $p$ of $C_{i}$ is positive (resp. negative).
We note that if the assumption $\kappa_{2}(p)=0$ is removed, the conclusion of Theorem 1.2 does not hold in general (Example 4.7).
@@ Acknowledgments
The author thanks Kentaro Saji and Runa Shimada for fruitful discussions and valuable comments. The author is also grateful to the referees for careful reading and constructive comments, which improved the manuscript.
@@ 2.1. Surfaces in Euclidean 3-space
We here review some fundamental concepts from differential geometry of immersed surfaces in Euclidean $3$-space $\mathbb{R}^{3}$. For details, see [3, 7, 18]. Let $U\subset\mathbb{R}^{2}$ be a domain. Let $f\colon U\to\mathbb{R}^{3}$ be an immersion and $\nu$ its unit normal vector field defined by
$\nu(u,v)=\frac{f_{u}\times f_{v}}{|f_{u}\times f_{v}|}(u,v),$
where $\times$ is the cross product of $\mathbb{R}^{3}$, $(\cdot)_{u}=\partial(\cdot)/\partial u$ and $(\cdot)_{v}=\partial(\cdot)/\partial v$. Then we call the image of $f$ or $f$ itself a (regular) surface. For a surface $f\colon U\to\mathbb{R}$, we define
(2.1) $E$ $=\left\langle{f_{u}},{f_{u}}\right\rangle,$ $F$ $=\left\langle{f_{u}},{f_{v}}\right\rangle,$ $G$ $=\left\langle{f_{v}},{f_{v}}\right\rangle,$
$L$ $=-\left\langle{f_{u}},{\nu_{u}}\right\rangle,$ $M$ $=-\left\langle{f_{u}},{\nu_{v}}\right\rangle=-\left\langle{f_{v}},{\nu_{u}}\right\rangle,$ $N$ $=-\left\langle{f_{v}},{\nu_{v}}\right\rangle,$
where $\left\langle{\cdot},{\cdot}\right\rangle$ is the Euclidean inner product. The functions $E,F,G$ are called the coefficients of the first fundamental form of $f$, and $L,M,N$ are the coefficients of the second fundamental form of $f$. Note that $EG-F^{2}>0$ holds on $U$.
Let $\widehat{\operatorname{I}}$ and $\widehat{\operatorname{II}}$ be matrices given by
$\widehat{\operatorname{I}}=\begin{pmatrix}E&F\\ F&G\end{pmatrix},\quad\widehat{\operatorname{II}}=\begin{pmatrix}L&M\\ M&N\end{pmatrix}$.
Then $\widehat{\operatorname{I}}$ is invertible. We define the Weingarten matrix $W$ as
$W=\widehat{\operatorname{I}}^{-1}\widehat{\operatorname{II}}=\frac{1}{EG-F^{2}}\begin{pmatrix}GL-FM&GM-FN\\ EM-FL&EN-FM\end{pmatrix}$.
The Gaussian curvature $K$ and mean curvature $H$ of $f$ are defined by
$K=\det{W},\quad H=\frac{1}{2}\operatorname{tr}W.$
Moreover, the eigenvalues $\kappa_{1},\kappa_{2}$ of $W$ are called the principal curvatures of $f$. A point $p\in U$ is called an umbilic point of $f$ if $\kappa_{1}(p)=\kappa_{2}(p)$.
Let $f\colon U\to\mathbb{R}^{3}$ be a surface and $\nu$ its unit normal vector field of $f$. Assume that $p\in U$ is not an umbilic point of $f$. Then there exist a sufficiently small neighborhood $\tilde{U}\subset U$ of $p$ and local coordinate system $(u,v)$ such that $F=M=0$ on $\tilde{U}$ (cf. [3, page 188]). We call such a coordinate system $(u,v)$ a curvature line coordinate system on $\tilde{U}$.
In the following, we assume that a surface $f\colon U\to\mathbb{R}^{3}$ has no umbilic point and parametrized by a curvature line coordinate system $(u,v)$. We define $\bm{e}_{1},\bm{e}_{2}\colon U\to\mathbb{R}^{3}$ by
(2.2) $\bm{e}_{1}=\frac{f_{u}}{|f_{u}|},\quad\bm{e}_{2}=\frac{f_{v}}{|f_{v}|}.$
Then replacing $\nu$ with $\nu$ if necessary it holds that
$\left\langle{\bm{e}_{i}},{\bm{e}_{j}}\right\rangle=\delta_{ij},\quad\bm{e}_{1}\times\bm{e}_{2}=\nu,$
where $\delta_{ij}$ is the Kronecker delta. The Gauss and Weingarten formulas are as follows:
@@
where $\kappa_{1}=L/E$ and $\kappa_{2}=N/G$. Moreover, Codazzi equations can be written as
@@ 2.2. Fronts
To investigate parallel surfaces below, we here recall the class of surfaces with certain singularities called fronts ([1, 25]). Let $f\colon U\to\mathbb{R}^{3}$ be a $C^{\infty}$ map, where $U\subset\mathbb{R}^{2}$ is an open set. Then $f$ is said to be a front if there exists a $C^{\infty}$ map $\nu\colon U\to\mathbb{S}^{2}$ such that
  1. (1)
@@
Let $f\colon U\to\mathbb{R}^{3}$ be a front with a unit normal vector field $\nu$. Then a point $p\in U$ is said to be a singular point of $f$ if $f$ is not an immersion at $p$, that is, $\operatorname{rank}df_{p}<2$. We denote the set of singular points of $f$ by $S(f)$. We define a function $\lambda\colon U\to\mathbb{R}$ by
@@
We then call $\lambda$ the signed area density function of $f$. It is clear that $S(f)=\lambda^{-1}(0)$ holds. Moreover, a function $\tilde{\lambda}$ which is non-zero proportional to $\lambda$ is called an identifier of singularities of $f$. A singular point $p\in S(f)$ of $f$ is said to be of non-degenerate if $(\tilde{\lambda}_{u},\tilde{\lambda}_{v})(p)\neq(0,0)$. Otherwise, it is said to be of degenerate. For a non-degenerate singular point $p\in S(f)$, there exist a neighborhood $V(\subset U)$ of $p$ and a regular curve $\gamma\colon(-\varepsilon,\varepsilon)\to V$ with $\gamma(0)=p$ such that $\tilde{\lambda}(\gamma(t))=0$ for $t\in(-\varepsilon,\varepsilon)$. This curve $\gamma$ is called a singular curve for $f$ passing through $p=\gamma(0)$. On the other hand, for a singular point $p\in S(f)$ satisfying $\operatorname{rank}df_{p}=1$, there exists a non-zero vector field $\eta$ on $V(\subset U)$ such that $\eta_{p}$ generates the kernel of $df_{p}$. This vector field $\eta$ is called a null vector field for $f$.
@@
Let $f_{i}\colon(U_{i},p_{i})\to\mathbb{R}^{3}$ ($i=1,2$) be $C^{\infty}$ map germs. Then $f_{1}$ and $f_{2}$ are $\mathcal{A}$-equivalent if there exist diffeomorphism germs $\varphi\colon(U_{1},p_{1})\to(U_{2},p_{2})$ and $\Phi\colon(\mathbb{R}^{3},f_{1}(p_{1}))\to(\mathbb{R}^{3},f_{2}(p_{2}))$ such that $\Phi\circ f_{1}=f_{2}\circ\varphi$.
@@
    $f$ is a cuspidal edge at $p(\in U)$ if $f$ is $\mathcal{A}$-equivalent to the germ $(u,v)\mapsto(u,v^{2},v^{3})$ at the origin;
@@
    $f$ is a swallowtail at $p$ if $f$ is $\mathcal{A}$-equivalent to the germ $(u,v)\mapsto(u,4v^{3}+2uv,3v^{4}+uv^{2})$ at the origin;
@@
    $f$ is a cuspidal butterfly at $p$ if $f$ is $\mathcal{A}$-equivalent to the germ $(u,v)\mapsto(u,5v^{4}+2uv,4v^{5}+uv^{2})$ at the origin;
@@
    $f$ is a cuspidal lips at $p$ if $f$ is $\mathcal{A}$-equivalent to the germ $(u,v)\mapsto(u,2v^{3}+u^{2}v,3v^{4}+u^{2}v^{2})$ at the origin;
@@
    $f$ is a cuspidal beaks at $p$ if $f$ is $\mathcal{A}$-equivalent to the germ $(u,v)\mapsto(u,2v^{3}-u^{2}v,3v^{4}-u^{2}v^{2})$ at the origin.